KTeV Shift Summary, Evening of 24 Dec 1999

Tony Barker (tonyb@cuhep.Colorado.EDU)
Fri, 24 Dec 1999 23:03:01 -0700 (MST)

Shift Summary for Evening Shift, 24 December 1999

On Shift: A. Alavi-Harati, T. Barker

Hours of beam delivered at all intensities: 7.5
Hours of beam delivered at nominal intensity: 7.5
Hours of E799 data (special runs not included): 7.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Miscellaneous KFRENDS and lost accelerator
spills; nothing major.

========================================================================
Runs this shift:

Run 15282 - 7.2M events - ENDED - BEAM_799
Run 15283 - 15.6M events - ONGOING - BEAM_799

Tapes brought to FCC this shift: UP(A,B,C)247-249, UPS083,
UPT020, UPE017, UPF062-063
UPK033, UPH045, UPM084, UPX017

========================================================================
Accesses during shift: none
========================================================================
Overview of shift:
Beam returns following D4 quench at 15:40,
followed by generally smooth running at about 9E12.
CsI constants updated after Run 15282 ended.
About 40 UPA,UPB,UPC tapes were labelled.
Flying reindeer spotted, heading south.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15282 248 15:40 Beam returns (was off since 13:57)
285 16:26 KTEVANA errors; run KFREND.
297 16:43 Tapes full; run ends.

16:50 Update CsI constants

15283 001 16:57 New run begins on new tapes;
run starts smoothly.
188 21:10 Trigger hung; KFREND to fix.
246 22:27 Trigger hung again; KFREND to fix.
264 22:50 KTEVANA errors; run KFREND;
followed by pipeline errors; reset pipeline.
273 23:07 Trigger hung again; KFREND to fix.
316 00:00 Hand off to Christmas OWL shift.
========================================================================
Follow up Notes on Previous Shift

- page Ed when a decently-sized run ends / tapes fill up.
---- Done at end of run 15282; CsI constants updated.

- someone should be showing up to repair the heater in the gas shed. Dave
Northacker is in charge of getting that done.
---- On-call electrician is here to see if anything can be done.
Apparently not without building electrical diagrams, which
are not at hand. Dave N. may look for them tomorrow, but
chances are nothing can be done until next week. Meanwhile
the gas 'shed' will be colder than a (well, you get the idea).

========================================================================
Items to Follow up on Next Shift

- see what BobT says about the CsI "heater" temperature.